A firm has the long-run cost function C(Q) = 4Q2 + 64.In the long run, it will supply a positive amount of output, so long as the price is greater than

a.

$64.

b.

$72.

c.

$16.

d.

$32.

e.

$37.
Answer #1

In long-run, firm will supply a positive amount of output as long as price is greater than long-run average total cost.

The given total cost function is as follows -

C(q) = 4q2 + 64

Calculate average total cost -

ATC = C/q

       = (4q2 + 64)/q

       = 4q + (64/q)

In order to calculate value of q we will first take the derivative of q and then equate it with zero.

Taking derivative of ATC with respect to q -

dATC/dq = d[4q + (64/q)]/dq = 4 - 64/q2

Equating derivative of ATC to find value of q -

4 - 64/q2 = 0

64/q2 = 4

q = 4

Calculating value of ATC -

ATC = 4q + 64/q = 4*4 + 64/4 = 16 + 16 = 32

Thus, ATC is $32.

As stated above, in long-run, firm will supply a positive amount of output as long as price is greater than long-run average total cost.

With long-run average total cost being $32, firm will supply a positive amount of output as long as price is greater than $32.

Thus, the correct answer is option (d).